The measures of the angles in a triangle are in the extended ratio of 3:4:5. What is the smallest angle of the triangle?

Answer by Cromlix(4381)

Add all the ratios
3+4+5 = 12
180 degrees in a triangle.
3/12 x 180 = 45
4/12 x 180 = 60
5/12 x 180 = 75
Total = 180
So, smallest angle = 45 degrees.
Hope this helps.
:-)